# Compactly supported $\mathbb{A}^{1}$-Euler characteristics of symmetric powers of cellular varieties

### v1

@@ 5.1. Grassmannians
Brazelton, McKean and Pauli [BMP23, Theorem 8.4] computed the $\mathbb{A}^{1}$-Euler characteristic of Grassmannians over a field $k$ which admits a real embedding $k\rightarrow\mathbb{R}$ by using a theorem of Bachmann and Wickelgren [BW23, Theorem 5.11]. We give a purely combinatorial proof which works over any field of characteristic zero, so without needing the condition that the field admits a real embedding. We will use Losanitsch’s triangle, OEIS-sequence A034851, which is a summand of Pascal’s triangle. It is a well-known combinatorial object constructed for example by Cigler [Cig17, Section 3]. The $d$-th entry in the $r$-th row is denoted by $e(d,r)$ and we define $o(d,r)=\binom{r}{d}-e(d,r)$. The numbers $e(d,r)$ and $o(d,r)$ satisfy the following recurrence relations:
@@
(1) $\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r))=e(d,r)\langle 1\rangle+o(d,r)\langle-1\rangle,$
@@
$\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r))=\chi^{\mathrm{mot}}(\operatorname{Gr}(d-1,r-1))+\langle(-1)^{d}\rangle\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r-1)),$
@@
$\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r))$ $=\chi^{\mathrm{mot}}(\operatorname{Gr}(d-1,r-1))+\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r-1)),$
@@
$\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r))$ $=\chi^{\mathrm{mot}}(\operatorname{Gr}(d-1,r-1))+\langle-1\rangle\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r-1)),$
@@
$\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r)^{(n)})=\sum_{i=0}^{n}\binom{e(d,r)+i-1}{i}\binom{o(d,r)+n-i-1}{n-i}\langle(-1)^{n-i}\rangle$
@@
Since Grassmannians are linear, they are $\operatorname{K}_{0}$-étale linear, so apply Theorem 4.10 to Theorem 5.2 to obtain
$\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r)^{(n)})$ $=a_{n}(\chi^{\mathrm{mot}}(\operatorname{Gr}(d,r)))$
@@
$\sum_{t=0}^{\infty}\chi^{mot}(\mathrm{Gr}(d,r)^{(n)})t^{n}=(1-t)^{-e(d,r)}(1-\langle-1\rangle t)^{-o(d,r)}\in\operatorname{GW}(k)[[t]].$
@@
In this subsection, we use the techniques from Section 4 to show a large class of del Pezzo surfaces are symmetrisable. These are a class of surfaces of arithmetic interest, famous for the fact that over the complex numbers they contain a finite number of lines lying on them. This number is well known from enumerative geometry. For example, smooth projective cubic surfaces are del Pezzo surfaces of degree $3$, which contain exactly $27$ lines over the complex numbers. Quadratically enriched counts of the number of lines on del Pezzo surfaces were achieved by Kass-Wickelgren [KW21, Theorem 2] in degree $3$, by Darwin [Dar22, Theorem 1.2] in degree $4$, and these were generalised to the degree $\geq 3$ case in [KLSW23b] and [KLSW23a] by Kass, Levine, Solomon and Wickelgren.
@@
A del Pezzo surface is a smooth projective variety of dimension $2$ whose anticanonical bundle is ample. The degree of a del Pezzo surface is the self intersection number of the anticanonical class. An exceptional curve on a del Pezzo surface is a curve with self intersection number $-1$.
Theorem 5.6 (Demazure, Théorème 1 of [Dem80]).
Let $\overline{k}$ be an algebraically closed field and let $X/\overline{k}$ be a del Pezzo surface of degree $d$. Then $1\leq d\leq 9$ and either:
@@
    $X$ is isomorphic to the blow up of $\mathbb{P}^{2}$ at $9-d$ points in general position.
While del Pezzo surfaces can be arithmetically very complicated, once we base change to the algebraic closure of our field, they are linear, so we would expect large classes of del Pezzo surfaces to be symmetrisable.
@@
This proceeds by checking on a case by case basis that the conditions we have already established for $X$ to be symmetrisable hold in our cases. We appeal to [Poo17, Section 9.4], and we sketch the main results here.
Suppose $X$ is a del Pezzo surface of degree $9$. Then $X$ is an even dimensional Severi Brauer variety, and since $X(k)\neq\emptyset$, we see $X\cong\mathbb{P}^{2}$, so $X$ is $\operatorname{K}_{0}$-étale linear and so symmetrisable by Theorem 4.10.
@@
Suppose $X$ is degree $7$. Then Proposition 9.4.17 of [Poo17] tells us that $X$ is isomorphic to the blow up of $\mathbb{P}^{2}$ at a closed subscheme isomorphic to a finite étale algebra of degree $2$, so the result holds by Corollary 4.13.
Suppose $X$ has degree $5$. Then by [Poo17, Proposition 9.4.20], if there exists a $k$-point of $X$ lying on none of the exceptional curves in $X$, we may obtain $X$ through blowing up and and blowing down $\mathbb{P}^{2}$. By [Poo17, Theorem 9.4.29], this is always the case in characteristic 00. Therefore, $X$ is symmetrisable.
Suppose $X$ is degree $6$, and let $x\in X(k)$. Using [Poo17, Proposition 9.4.20], if $x$ lies on exactly one exceptional curve $E_{i}$, then we may blow down $X$ to obtain a del Pezzo surface of degree $7$. Therefore, $X$ is symmetrisable by the degree $7$ case and Corollary 4.13. If $x$ lies on an intersection of exceptional curves, then we may blow down two other exceptional curves to express $X$ obtain a del Pezzo surface of degree $8$, so $X$ is symmetrisable by Corollary 4.13 and the degree $8$ case. Finally, if $x$ does not lie on any exceptional curves, we blow up $X$ at $x$ to obtain a del Pezzo surface of degree $5$, so $X$ is symmetrisable by the degree $5$ case and Corollary 4.13. ∎
If $X$ is a del Pezzo surface of degree $5$ or $7$, the condition that $X(k)\neq\emptyset$ is immediate. If $X$ is a del Pezzo surface of degree $9$, then it is a Severi–Brauer surface so it is symmetrisable by Corollary 4.19. We also see del Pezzo surfaces of degree $6$ which are given by blow ups of Severi–Brauer surfaces at a point are also symmetrisable by Corollary 4.13, so the above also holds for certain del Pezzo surfaces with no $k$-point. We can also use our results to show the following.
@@
is $\operatorname{K}_{0}$-étale linear. Note that $Y$ contains $2$ skew lines defined over $k$: namely, the lines $L_{1}:\{x=t,y=-z\}$ and $L_{2}:\{x=-t,y=z\}$. Therefore we may blow $Y$ down at these $2$ lines to obtain a del Pezzo surface of degree $5$. This del Pezzo surface will have a $k$-point, since the skew lines are defined over $k$, so is symmetrisable by Theorem 5.7, and therefore $Y$ is symmetrisable by Corollary 4.13.
@@
[…]
@@
$\chi^{\mathrm{mot}}(X)=2\langle 1\rangle+4\langle-1\rangle+\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle.$
@@
By the blow up formula for $\chi^{\mathrm{mot}}$, we see that
$\chi^{\mathrm{mot}}(X)=\chi^{\mathrm{mot}}(\mathbb{P}^{2})+\chi^{\mathrm{mot}}(E)-\chi^{\mathrm{mot}}(Y),$
where $E$ is the exceptional divisor of the blow up. The exceptional divisor is a $\mathbb{P}^{1}$ bundle over $Y$, so $\chi^{\mathrm{mot}}(E)=\chi^{\mathrm{mot}}(\mathbb{P}^{1})\cdot\chi^{\mathrm{mot}}(Y)$. Note that $\chi^{\mathrm{mot}}(\mathbb{P}^{1})=\mathbb{H}$, so $\chi^{\mathrm{mot}}(E)-\chi^{\mathrm{mot}}(Y)=\langle-1\rangle\cdot\chi^{\mathrm{mot}}(Y)$. By [Hoy14, Proposition 5.2], we see that if $L/k$ is a finite field extension, then $\chi^{\mathrm{mot}}(\operatorname{Spec}(L))=\mathrm{Tr}_{L/k}$, where $[\mathrm{Tr}_{L/k}]$ is the trace form on $L$. If $L=k(\sqrt{\alpha})$, then computing the trace form in the basis $1,\sqrt{\alpha}$ gives $\chi^{\mathrm{mot}}(\operatorname{Spec}(L))=\langle 2\rangle+\langle 2\alpha\rangle=\langle 1\rangle+\langle\alpha\rangle$. Additivity of $\chi^{\mathrm{mot}}$ implies
$\chi^{\mathrm{mot}}(Y)=3\langle 1\rangle+\langle\alpha\rangle+\langle\beta\rangle+\langle\gamma\rangle.$
Finally, $\chi^{\mathrm{mot}}(\mathbb{P}^{2})=2\langle 1\rangle+\langle-1\rangle$. Put together, this gives
$\chi^{\mathrm{mot}}(X)=2\langle 1\rangle+4\langle-1\rangle+\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle.$
@@
$\chi^{\mathrm{mot}}(X^{(3)})$ $=60\mathbb{H}+14\langle-1\rangle+8\left(\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle\right)$
$+2\left(\langle-\alpha\beta\rangle+\langle-\alpha\gamma\rangle+\langle-\beta\gamma\rangle\right)+\langle-\alpha\beta\gamma\rangle+t_{\alpha\beta}+t_{\beta\gamma}+t_{\alpha\gamma}.$
@@
Since $X$ is symmetrisable, we see that $\chi^{\mathrm{mot}}(X^{(3)})=a_{3}(\chi^{\mathrm{mot}}(X))$. We may compute $a_{3}(\chi^{\mathrm{mot}}(X))$, using the additive formula for the $a_{n}$s, to obtain
$\chi^{\mathrm{mot}}(X^{(3)})$ $=a_{3}(2\langle 1\rangle+4\langle-1\rangle)$
$+a_{2}(2\langle 1\rangle+4\langle-1\rangle)\cdot\left(\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle\right)$
$+(2\langle 1\rangle+4\langle-1\rangle)\cdot a_{2}(\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle)$
$+a_{3}(\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle).$
Write $\phi=\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle$ to ease notation. Standard computations utilising [PP23, Lemmas 3.18 and 3.19] give us
$a_{3}(2\langle 1\rangle+4\langle-1\rangle)$ $=24\mathbb{H}+8\langle-1\rangle$
$a_{2}(2\langle 1\rangle+4\langle-1\rangle)\phi$ $=24\mathbb{H}+5\phi$
$(2\langle 1\rangle+4\langle-1\rangle)\cdot a_{2}(\phi)$ $=12\mathbb{H}+6\langle-1\rangle+2\left(\langle-\alpha\beta\rangle+\langle-\beta\gamma\rangle+\langle-\alpha\gamma\rangle\right)$
$a_{3}(\phi)$ $=3\phi+\langle-\alpha\beta\gamma\rangle+t_{\alpha\beta}+t_{\beta\gamma}+t_{\alpha\gamma},$
and putting this together gives
$\chi^{\mathrm{mot}}(X^{(3)})$ $=60\mathbb{H}+14\langle-1\rangle+8\left(\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle\right)$
$+2\left(\langle-\alpha\beta\rangle+\langle-\alpha\gamma\rangle+\langle-\beta\gamma\rangle\right)+\langle-\alpha\beta\gamma\rangle+t_{\alpha\beta}+t_{\beta\gamma}+t_{\alpha\gamma}.$

### v2

@@ 5.1. Grassmannians
It is well known that Grassmannians are linear varieties in the sense of Definition 3.8, see for example [Jos01, Example 2.2], and are therefore symmetrisable by Theorem 4.10. In this subsection, we compute the Euler characteristic of symmetric powers of Grassmannians. Brazelton, McKean and Pauli [BMP23, Theorem 8.4] computed the $\mathbb{A}^{1}$-Euler characteristic of Grassmannians over a field $k$ which admits a real embedding $k\rightarrow\mathbb{R}$ by using a theorem of Bachmann and Wickelgren [BW23, Theorem 5.11]. We give a purely combinatorial proof which works over any field of characteristic $\neq 2$, so without needing the condition that the field admits a real embedding. We will use Losanitsch’s triangle, OEIS-sequence A034851, which is a summand of Pascal’s triangle. It is a well-known combinatorial object constructed for example by Cigler [Cig17, Section 3]. The $d$-th entry in the $r$-th row is denoted by $e(d,r)$ and we define $o(d,r)=\binom{r}{d}-e(d,r)$. The numbers $e(d,r)$ and $o(d,r)$ satisfy the following recurrence relations:
@@
(1) $\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r))=e(d,r)\langle 1\rangle+o(d,r)\langle-1\rangle,$
@@
$\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r))=\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d-1,r-1))+\langle(-1)^{d}\rangle\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r-1)),$
@@
$\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r))$ $=\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d-1,r-1))+\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r-1)),$
@@
$\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r))$ $=\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d-1,r-1))+\langle-1\rangle\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r-1)),$
@@
$\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r)^{(n)})=\sum_{i=0}^{n}\binom{e(d,r)+i-1}{i}\binom{o(d,r)+n-i-1}{n-i}\langle(-1)^{n-i}\rangle.$
@@
Since Grassmannians are linear, they are $\mathrm{K}_{0}$-étale linear, so apply Theorem 4.10 to Theorem 5.1 to obtain
$\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r)^{(n)})$ $=a_{n}(\chi_{c}^{\mathrm{mot}}(\operatorname{Gr}(d,r)))$
@@
$\sum_{n=0}^{\infty}\chi_{c}^{\mathrm{mot}}(\mathrm{Gr}(d,r)^{(n)})t^{n}=(1-t)^{-e(d,r)}(1-\langle-1\rangle t)^{-o(d,r)}\in\operatorname{GW}(k)[[t]].$
@@
Corollary 5.4.
Take $k=\mathbb{R}$ to be our base field, and consider $\mathrm{Gr}(d,r)/k$. For $T$ a CW-complex, write $e_{c}(T)$ to mean the compactly supported Euler characteristic of $T$. Then compactly supported Euler characteristics of the complex and real points of $\mathrm{Gr}(d,r)^{(n)}$ respectively fit into the following power series in $\mathbb{Z}[[t]]$
$\sum_{n=0}^{\infty}e_{c}(\mathrm{Gr}(d,r)^{(n)})(\mathbb{C}))^{n}$ $=(1-t)^{n\choose d}$,
$\sum_{n=0}^{\infty}e(\mathrm{Gr}(d,r)^{(n)}(\mathbb{R}))t^{n}$ $=(1-t)^{-e(d,r)}(1+t)^{-o(d,r)}.$
Proof.
The first result follows by applying the homomorphism $\mathrm{rank}:\operatorname{GW}(\mathbb{R})\to\mathbb{Z}$ and using [Lev20, Remark 2.3.1] to see that $\mathrm{rank}(\chi_{c}^{\mathrm{mot}}(X))=e(X(\mathbb{C}))$ (see [PPar, Theorem 2.14] for the extension of this result to the compactly supported case). The result about the real points follows by instead applying $\mathrm{sign}:\operatorname{GW}(\mathbb{R})\to\mathbb{Z}$ and using [Lev20, Remark 2.3.1] which shows that $\mathrm{sign}(\chi_{c}^{\mathrm{mot}}(X))=e_{c}(X(\mathbb{R}))$ (see also [PPar, Theorem 2.20] for the extension of this result to the compactly supported case). ∎
@@
In this subsection, we use the techniques from Section 4 to show a large class of del Pezzo surfaces are symmetrisable. These are a class of surfaces of arithmetic interest, famous for the fact that over the complex numbers they contain a finite number of exceptional curves lying on them. This number is well known from enumerative geometry, for example: smooth projective cubic surfaces are del Pezzo surfaces of degree $3$, the exceptional curves are precisely the lines on the surface, and it is well known that exactly $27$ lines lie on the surface over the complex numbers. Quadratically enriched counts of the number of exceptional curves on del Pezzo surfaces were achieved by Kass-Wickelgren [KW21, Theorem 2] in degree $3$, by Darwin [Dar22, Theorem 1.2] in degree $4$, and these were generalised to the degree $\geq 3$ case in [KLSW23b] and [KLSW23a] by Kass, Levine, Solomon and Wickelgren.
@@
A del Pezzo surface is a smooth projective variety of dimension $2$ whose anticanonical bundle is ample. The degree of a del Pezzo surface is the self intersection number of the anticanonical class. An exceptional curve on a del Pezzo surface is a curve with self intersection number 1-1.
The following classification of del Pezzo surfaces is due to Manin over algebraically closed fields, which we can weaken to separably closed fields by a result of Coombes.
Theorem 5.6 (Theorem 24.4 of [Man86], Theorem 1 of [Coo88]).
Let $\overline{k}$ be a separably closed field and let $X/\overline{k}$ be a del Pezzo surface of degree $d$. Then $1\leq d\leq 9$ and either:
@@
    $X$ is isomorphic to the blow up of $\mathbb{P}^{2}$ at 9d9-d points in general position.
While del Pezzo surfaces can be arithmetically very complicated, once we base change to the separable closure of our field, they are linear, so we would expect large classes of del Pezzo surfaces to be symmetrisable.
@@
This proceeds by checking on a case by case basis that the conditions we have already established for $X$ to be symmetrisable for such del Pezzo surfaces. We appeal to [Poo17, Section 9.4], and we sketch the main results here.
Suppose $X$ is a del Pezzo surface of degree $9$. Then $X$ is an even dimensional Severi Brauer variety, and since $X(k)\neq\emptyset$, we see $X\cong\mathbb{P}^{2}$, so $X$ is $\mathrm{K}_{0}$-étale linear and so symmetrisable by Theorem 4.10.
@@
Suppose $X$ is degree $7$. Then by [Poo17, Proposition 9.4.17], $X$ is isomorphic to the blow up of $\mathbb{P}^{2}$ at a closed subscheme isomorphic to a finite étale algebra of degree $2$, so the result holds by Corollary 4.13. Note that this Proposition does not require $X$ to have a $k$ point.
Suppose $X$ has degree $5$. Then by [Poo17, Proposition 9.4.20], if there exists a $k$-point of $X$ lying on none of the exceptional curves in $X$, we may obtain $X$ through blowing up and and blowing down $\mathbb{P}^{2}$. By [Poo17, Theorem 9.4.29], this is always the case unless $k=\mathbb{F}_{2},\mathbb{F}_{4}$ or $\mathbb{F}_{3}$. Therefore $X$ is symmetrisable under our assumptions on $k$ by Corollary 4.13 unless $k=\mathbb{F}_{3}$, since we are assuming $k$ has characteristic not $2$. It only remains to tackle the case that $k=\mathbb{F}_{3}$ and all $k$ points of $X$ lie on the exceptional curves in $X$. In this case $X$ is isomorphic to $\mathbb{P}^{2}_{k}$ blown up at $4$ points by [Poo17, Theorem 9.4.29], which gives the result by Corollary 4.13.
Suppose $X$ is degree $6$, and let $x\in X(k)$. Using [Poo17, Proposition 9.4.20], if $x$ lies on exactly one exceptional curve $E_{i}$, then we may blow down $X$ to obtain a del Pezzo surface of degree $7$. Therefore, $X$ is symmetrisable by the degree $7$ case and Corollary 4.13. If $x$ lies on an intersection of exceptional curves, then we may blow down two other exceptional curves to obtain a del Pezzo surface of degree $8$, so $X$ is symmetrisable by Corollary 4.13 and the degree $8$ case. Finally, if $x$ does not lie on any exceptional curves, we blow up $X$ at $x$ to obtain a del Pezzo surface of degree $5$, so $X$ is symmetrisable by the degree $5$ case and Corollary 4.13. ∎
Remark 5.8.
Even without the assumption that $X(k)\neq\emptyset$, if $X$ is a del Pezzo surface of degree $5$ or $7$, then the above work and [Poo17, Theorem 9.4.29] tells us that $X$ is birational to $\mathbb{P}^{2}$. If $k$ is infinite, the $k$ points of $\mathbb{P}^{2}$ are Zariski dense, and so we automatically have that $X(k)\neq\emptyset$. If $X$ is a del Pezzo surface of degree $9$, then it is a Severi–Brauer surface. We also see del Pezzo surfaces of degree $6$ which are given by blow ups of Severi–Brauer surfaces at a point are also symmetrisable by Corollary 4.13, so the above also holds for certain del Pezzo surfaces with no $k$-point.
The reason we restrict to the degree $\geq 5$ case above is that for del Pezzo surfaces $X$ of degree $\geq 5$ such that $X(k)\neq\emptyset$, we have that $X$ is birational to $\mathbb{P}^{2}$, by the steps taken in the proof above. We can therefore obtain all such del Pezzo surfaces by iterating blow-ups and blow-downs, which do not change symmetrisability by Corollary 4.13.
When $X$ has degree $\leq 4$, this is no longer true: there exist del Pezzo surfaces of degree $\leq 4$ which are not birational to $\mathbb{P}^{2}$. However we can obtain similar results using Corollary 4.13 for any del Pezzo surface of degree $\leq 4$ which can be obtained via iterating blow ups and blow downs of $\mathbb{P}^{2}$.
We can also use our results to show the following.
@@
is $\mathrm{K}_{0}$-étale linear. Note that $Y$ contains $2$ skew lines defined over $k$: namely, the lines $L_{1}:\{x=t,y=-z\}$ and $L_{2}:\{x=-t,y=z\}$. Therefore we may blow $Y$ down at these $2$ lines to obtain a del Pezzo surface of degree $5$. This del Pezzo surface will have a $k$-point, since the skew lines are defined over $k$, so is symmetrisable by Theorem 5.7, and therefore $Y$ is symmetrisable by Corollary 4.13.
@@
for some $a_{1},a_{2},a_{3}\in k^{\times}$. Therefore, $X$ and $Y$ become isomorphic once we base change to the field $L:=k(\sqrt[3]{a_{1}},\sqrt[3]{a_{2}},\sqrt[3]{a_{3}})$. Note that $L:k]=3^{i}$ where $i\in\{0,1,2,3\}$. In particular, we can apply Theorem 4.16 to obtain the result. ∎
@@
$\chi_{c}^{\mathrm{mot}}(X)=2\langle 1\rangle+4\langle-1\rangle+\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle.$
@@
By the blow up formula for $\chi_{c}^{\mathrm{mot}}$, we see that
$\chi_{c}^{\mathrm{mot}}(X)=\chi_{c}^{\mathrm{mot}}(\mathbb{P}^{2})+\chi_{c}^{\mathrm{mot}}(E)-\chi_{c}^{\mathrm{mot}}(Y),$
where $E$ is the exceptional divisor of the blow up. The exceptional divisor is a $\mathbb{P}^{1}$ bundle over $Y$, so $\chi_{c}^{\mathrm{mot}}(E)=\chi_{c}^{\mathrm{mot}}(\mathbb{P}^{1})\cdot\chi_{c}^{\mathrm{mot}}(Y)$. Note that $\chi_{c}^{\mathrm{mot}}(\mathbb{P}^{1})=\mathbb{H}$, so $\chi_{c}^{\mathrm{mot}}(E)-\chi_{c}^{\mathrm{mot}}(Y)=\langle-1\rangle\cdot\chi_{c}^{\mathrm{mot}}(Y)$. By [Hoy14, Proposition 5.2], we see that if $L/k$ is a finite field extension, then $\chi_{c}^{\mathrm{mot}}(\operatorname{Spec}(L))=[\mathrm{Tr}_{L/k}]$, where $\mathrm{Tr}_{L/k}]$ is the trace form on $L$. If $L=k(\sqrt{\alpha})$, then computing the trace form in the basis $1,\sqrt{\alpha}$ gives $\chi_{c}^{\mathrm{mot}}(\operatorname{Spec}(L))=\langle 2\rangle+\langle 2\alpha\rangle$. Additivity of $\chi_{c}^{\mathrm{mot}}$ implies
$\chi_{c}^{\mathrm{mot}}(Y)=3\langle 2\rangle+\langle 2\alpha\rangle+\langle 2\beta\rangle+\langle 2\gamma\rangle=2\langle 1\rangle+\langle 2\rangle+\langle 2\alpha\rangle+\langle 2\beta\rangle+\langle 2\gamma\rangle$
Finally, $\chi_{c}^{\mathrm{mot}}(\mathbb{P}^{2})=2\langle 1\rangle+\langle-1\rangle$. Put together, this gives
$\chi_{c}^{\mathrm{mot}}(X)=2\mathbb{H}+\langle-1\rangle+\langle-2\rangle+\langle-2\alpha\rangle+\langle-2\beta\rangle+\langle-2\gamma\rangle.$
@@
$\chi_{c}^{\mathrm{mot}}(X^{(3)})$ $=60\mathbb{H}+11\langle-1\rangle+3\langle-2\rangle$
$+7\left(\langle-2\alpha\rangle+\langle-2\beta\rangle+\langle-2\gamma\rangle\right)+\left(\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle\right)$
$+\left(\langle 1\rangle+\langle 2\rangle\right)\left(\langle\alpha\beta\rangle+\langle\alpha\gamma\rangle+\langle\beta\gamma\rangle\right)$
$+\langle-2\alpha\beta\gamma\rangle+t_{\alpha\beta}+t_{\beta\gamma}+t_{\alpha\gamma}$
@@
Since $X$ is symmetrisable, we see that $\chi_{c}^{\mathrm{mot}}(X^{(3)})=a_{3}(\chi_{c}^{\mathrm{mot}}(X))$. We may compute $a_{3}(\chi_{c}^{\mathrm{mot}}(X))$, using the additive formula for the $a_{n}$s, to obtain
$\chi_{c}^{\mathrm{mot}}(X^{(3)})$ $=a_{3}(2\langle 1\rangle+3\langle-1\rangle+\langle-2\rangle)$
$+a_{2}(2\langle 1\rangle+3\langle-1\rangle+\langle-2\rangle)\cdot\left(\langle-2\alpha\rangle+\langle-2\beta\rangle+\langle-2\gamma\rangle\right)$
$+(2\langle 1\rangle+3\langle-1\rangle+\langle-2\rangle)\cdot a_{2}(\langle-2\alpha\rangle+\langle-2\beta\rangle+\langle-2\gamma\rangle)$
$+a_{3}(\langle-2\alpha\rangle+\langle-2\beta\rangle+\langle-2\gamma\rangle).$
Write $\phi=\langle-2\alpha\rangle+\langle-2\beta\rangle+\langle-2\gamma\rangle$ to ease notation. Standard computations utilising [PP25, Lemmas 3.20 and 3.21] give us
$a_{3}(2\mathbb{H}+\langle-1\rangle+\langle-2\rangle)$ $=24\mathbb{H}+8\langle-1\rangle$
$a_{2}(2\mathbb{H}+\langle-1\rangle+\langle-2\rangle)\phi$ $=24\mathbb{H}++4\phi+\langle 2\rangle\phi$
$(2\langle 1\rangle+3\langle-1\rangle+\langle-2\rangle)\cdot a_{2}(\phi)$ $=12\mathbb{H}+(\langle-1\rangle+\langle-2\rangle)\left(3\langle 1\rangle+\langle-\alpha\beta\rangle+\langle-\beta\gamma\rangle+\langle-\alpha\gamma\rangle\right)$
$a_{3}(\phi)$ $=3\phi+\langle-2\alpha\beta\gamma\rangle+t_{\alpha\beta}+t_{\beta\gamma}+t_{\alpha\gamma}.$
Compiling these terms and substituting our value back for $\phi$ gives
$\chi_{c}^{\mathrm{mot}}(X^{(3)})$ $=60\mathbb{H}+11\langle-1\rangle+3\langle-2\rangle$
$+7\left(\langle-2\alpha\rangle+\langle-2\beta\rangle+\langle-2\gamma\rangle\right)+\left(\langle-\alpha\rangle+\langle-\beta\rangle+\langle-\gamma\rangle\right)$
$+\left(\langle 1\rangle+\langle 2\rangle\right)\left(\langle\alpha\beta\rangle+\langle\alpha\gamma\rangle+\langle\beta\gamma\rangle\right)$
$+\langle-2\alpha\beta\gamma\rangle+t_{\alpha\beta}+t_{\beta\gamma}+t_{\alpha\gamma}.$
Remark 5.12.
It should be noted that Pepin-Lehalleur and Taelman have work in progress which computes $X^{(n)}$ for a very general class of surfaces $X$. Moreover, Section 8 of [BM25] to derives a Göttsche style formula for computing the compactly supported $\mathbb{A}^{1}$-Euler characteristic of Hilbert schemes of points of symmetrisable surfaces. This result is of particular interest if we can show that $\mathrm{K}3$ surfaces are symmetrisable, when $X$ is a $\mathrm{K}3$-surface, this formula would generalise [PPar, Corollary 8.18]. In characteristic 0 we see that $\mathrm{K}3$ surfaces are not in general geometrically stably rational (for example: the unramified Brauer group of a variety is a stable birational invariant, and we may find examples of $\mathrm{K}3$ surfaces whose unramified Brauer group is non-trivial). Therefore $\mathrm{K}3$ surfaces are not in general $\mathrm{K}_{0}$-étale linear by Theorem 3.3, and so the results of this paper do not apply to give such a formula. However, we may apply Theorem 8.7 of [BM25], to obtain a Göttsche formula for symmetrisable surface, and in particular, the results of our paper implies that the formula in Theorem 8.7 of [BM25] holds whenever $X/k$ is a del Pezzo surface of degree $\geq 5$ with $X(k)\neq\emptyset$.